OpenStudy (anonymous):

Calculus help? Given the derivative of f(x), f ‘(x) = x^3 + 2x^2 – 8x – 8, a) Use your calculator to determine one x value where f(x) has a relative extrema. Your value should be correct to 3 decimal places. b) Classify the x value as a relative maximum or minimum. Justify your answer in a complete sentence. why do they tell me to use a calculator? am I supposed to insert a number to make it equal 0??

OpenStudy (kc_kennylau):

Relative extrema are found when f'(x)=0

OpenStudy (kc_kennylau):

Then, if f''(x) is negative, it's a local maximum. If f''(x) is positive, it's a local minimum.

OpenStudy (kc_kennylau):

P.S. If f''(x) is zero, double-check your calculation xD

OpenStudy (anonymous):

so I need to factor it right?

OpenStudy (kc_kennylau):

I'm not sure, because normal schools won't ask you to factorize a cubic equation...

OpenStudy (kc_kennylau):

Well, it says "use a calculator", so I'm not sure how to do it.

OpenStudy (anonymous):

0=x^2(x+2) 8(x-1) is this correct?

OpenStudy (anonymous):

yeah thats whats confusing me?

OpenStudy (kc_kennylau):

@ganeshie8 @dumbcow ADMIN: @hartnn @satellite73 @.Sam. @amistre64 @Callisto

OpenStudy (kc_kennylau):

Please ban meghen_elodie.

OpenStudy (anonymous):

You can not factor it easily, you may check the graph of this function on https://www.desmos.com/calculator. In this graph, I see we have 3 points of x, one of them maybe 9/4. Other points you should estimate or solve by yourself. Hint: (x-9/4)(Ax^2+Bx+C), then find A,B,C easily

OpenStudy (kc_kennylau):

(Meghen: you're lucky, no admins are on now, but I've reported you, so don't worry)

OpenStudy (anonymous):

I am not sure how @linh412986 , kc can you see the points?

OpenStudy (kc_kennylau):

@linh412986 adding a full-stop has made the website unaccessable, and the x-intercept is definitely not 9/4

OpenStudy (kc_kennylau):

https://www.desmos.com/calculator/rqnvfbejsk

OpenStudy (kc_kennylau):

@VIbarguen1 where did you get this question? :)

OpenStudy (anonymous):

@VIbarguen1 The purpose of question a) is explained by @kc_kennylau , we must find the solution of f' = 0 to find the extrema points. In order to solve for f' = 0, we solve the cubic equation. But it is hard to solve it, instead of it, we estimate the solution of f'=0 by graphing on desmos. @kc_kennylau I made a mistake, one solution should be 5/2

OpenStudy (kc_kennylau):

@linh412986 neither is it 5/2.

OpenStudy (anonymous):

it's homework :(

OpenStudy (kc_kennylau):

@linh412986 that is just approximation. the zeros calculated by WolframAlpha are here: http://www.wolframalpha.com/input/?i=factorize+x%5E3%2B2x%5E2-8x-8

OpenStudy (loser66):

there are 2 ways to use calculator to find out the roots

OpenStudy (anonymous):

Yes, that's just an approximation from graph. But the point here is that requirement of using calculator to solve :)

OpenStudy (anonymous):

so kc you mean 2.494, -0.890, and 3.603 ?

OpenStudy (kc_kennylau):

https://en.wikipedia.org/wiki/Newton's_method

OpenStudy (anonymous):

lin, I don't know how

OpenStudy (anonymous):

@Loser66 . can you teach how to use calculator for this case?

OpenStudy (anonymous):

it doesn't show it right

OpenStudy (loser66):

1/ hit the key Y= ..... then type the expression and then hit the trace key to find the roots 2/use program key to solve the cubic

OpenStudy (kc_kennylau):

You can use the Newton's method: \[\Huge x_{n+1}=x_n-\frac{f(x_n)}{f'(x_n)}\](where \(x_{n+1}\) is the more accurate version of \(x_n\)) https://en.wikipedia.org/wiki/Newton's_method

OpenStudy (loser66):

you can hit the table key to see the table of the graph then up or down to see where y =0

OpenStudy (loser66):

by using program, I have x1 = -3.603875472 x2= -0.8900837358 x3= 2.493959270

OpenStudy (anonymous):

I can do the same on an online cal, like the one lin put up ys?

OpenStudy (kc_kennylau):

Is the Newton's Method allowed?

OpenStudy (anonymous):

thats what kc sent me with wolframalpha

OpenStudy (anonymous):

I don't know. @Loser66 is it?

OpenStudy (loser66):

I don't follow, sorry friends, just support you in finding the roots. Follow @kc_kennylau guiding, please

OpenStudy (kc_kennylau):

What has your school taught you? Maybe you can look from the previous examples in your book?

OpenStudy (anonymous):

There isn't any example quite like this one in the book. and I do online schooling so the book is the only thing I have to follow :(

OpenStudy (anonymous):

at least, I can get this problem done.

OpenStudy (anonymous):

trust me, it's been thanks to the help I've gotten on here that can I understand how to do half of the problems but some I still can't solve.

OpenStudy (loser66):

http://www.youtube.com/watch?v=bITQg_k7K0c

OpenStudy (anonymous):

thanks you guys, for all the help

OpenStudy (kc_kennylau):

no problem at all :)

OpenStudy (loser66):

http://www.youtube.com/watch?v=8MfMiLbunuY if you are allowed to use calculator to solve quadratic and cubic, you should have those program on yours. for your problem f"(x) = 0 when x = 1.097167541 and x = -2.430500874 so far, you have 5 candidates for critical points ( extrema and local max/min) to find value of x , you just take integral of f'(x) to find out the original function of f (x) then plug those values to find the values of f(x). Compare them to consider which is extreme max/min, which is local one . DAT SIT

OpenStudy (anonymous):

I did that and I got 2.494 to work. Am I right?

OpenStudy (loser66):

\[\int x^3 +2x^2-8x-8 dx = \frac{x^4}{4}+\frac{2x^3}{3}-4x^2-8x +C\]

OpenStudy (anonymous):

ok got it, thank you!!

OpenStudy (loser66):

for x = -3.604 f(-3.604) = -12.154 do the same with the leftover.

OpenStudy (anonymous):

ok I will

OpenStudy (loser66):

good luck

OpenStudy (anonymous):

thanks